Solve the proportion:
(x-1)/3=(x+5)/(x-1)
Thanks for helping me out!!

Answer by stanbon(75887)

Solve the proportion:
(x-1)/3=(x+5)/(x-1)
Cross-multiply to get:
3(x+5) = (x-1)^2
3x+15 = x^2-2x+1
x^2-5x-14 = 0
(x-7)(x+2)=0
x=7 or x=-2
=============
Cheers,
Stan H.